# Biharmonic nonlinear vector field equations in dimension 44

### v2

@@ 1 Introduction and main results
Due to the increasing interest for higher order models and their applications (e.g. in nonlinear elasticity [1] and mechanics [17]), it is also relevant to investigate ground state problems for the bilaplacian operator $\Delta^{2}$. This was done by Mederski and Siemienowski [16] for dimensions $d\geq 5$, and as an application they conceived a new biharmonic logarithmic Sobolev inequality (cf. [16, Theorem 1.4]).
The scope of the present paper is to study the limiting case of space dimension $4$. Our approach is based on the method introduced in [8], as well as on some tools borrowed from [16] (e.g. Pohozaev identity in [16, Theorem 1.2]). Following [16], we also extend the biharmonic logarithmic Sobolev inequality in [16, Theorem 1.4] to dimension $4$.
@@
where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$, $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ [^1] To simplify the notation, we write $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ instead of $g\in L_{loc}^{\infty}(\mathbb{R}^{m};\mathbb{R}^{m})$, and when considering other vector-valued maps belonging to some $L^{p}$ space, we shall in general omit the range., is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
@@
$G(u_{0})>0\ \text{for some}\ u_{0}.$ (6)
That is, the function $G$ has at $u=0$ a local (but not global) maximum equal to 0. We also point out that $G$ need not be differentiable at 0. For instance, we may have G(u)=|u|G(u)=-|u| near 0. The energy functional associated with system (2) is
@@
$\text{where }K(u)=\dfrac{1}{2}\int_{\mathbb{R}^{4}}\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int_{\mathbb{R}^{4}}\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (8)
$\text{ and }V(u)=\int_{\mathbb{R}^{4}}G(u(x))dx.$ (9)
@@
It is worth noting that $K(u)=\frac{1}{2}\int_{\mathbb{R}^{4}}|\Delta u|^{2}$, provided that $\nabla u\in D^{1,2}(\mathbb{R}^{4})$ (cf. Lemma 11). Thus, $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$. We shall also see in Lemma 1 below, that the class $\mathcal{C}$ is contained in $L^{\infty}(\mathbb{R}^{4})$. Due to this property, and in contrast with the case of the second order system (1), none asymptotic condition on $G$ is required to prove the existence of a ground state solution to (2).
@@
If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$.
@@
which clearly satisfies assumptions (3)-(6). Moreover, for this choice of $G$, one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Following [16], we derive in (12) below the biharmonic logarithmic Sobolev inequality in dimension $4$ (cf. [16, Theorem 1.4] for dimensions $d\geq 5$). It is a generalization for the bilaplacian operator of the classical logarithmic Sobolev inequality
@@
Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$. For any $u\in\mathcal{C}^{\prime}$ there holds
@@
with $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$, where $T>0$ is the minimum in (10). Moreover, the equality in (12) holds provided that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ and $\bar{u}$ is a ground state solution to (2). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
@@
Equality case in (12) is achieved for the normalized ground state solutions provided by Theorem 1.1. However, these solutions are not known explicitly, thus we neither know the exact value of the optimal constant in (12). By comparing it with the optimal constant (cf. [9]) of the classical logarithmic Sobolev inequality (11), we can at least obtain that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in the Appendix we gather some technical lemmas.
@@
$\mu(\{|u_{j}|>\varepsilon\})=1.$ (13)
It is clear that the norms $\|D^{2}u_{j}\|_{L^{2}(\mathbb{R}^{4})}$ are uniformly bounded. Therefore, due to the Sobolev imbedding of $D^{1,2}(\mathbb{R}^{4})$ in $L^{4}(\mathbb{R}^{4})$ (cf. [20, Theorem 1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
Next, we establish some properties of the class $\mathcal{C}$, and the sequence $\{u_{j}\}$.
@@
Let $v\in\mathcal{C}$, then $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, $\forall\alpha\in(0,1)$, and it holds $\lim_{|x|\to\infty}v(x)=0$. In addition, the sequence $\{u_{j}\}$ is equibounded as well as equicontinuous in $\mathbb{R}^{4}$, and satisfies
$\int_{\{|u_{j}|>\varepsilon\}}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (14)
@@
We have for every $q\geq 1$ the following estimates.
$\|\psi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\psi\|_{L^{2}(\mathbb{R}^{4})}(\mu(supp\psi))^{1/q},$ (15a)
holding for $\psi\in L^{1}_{loc}(\mathbb{R}^{4})$ such that $\nabla\psi\in L^{2}(\mathbb{R}^{4})$ and $\mu(supp\psi)<\infty$,
$\|\phi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{\frac{1}{q}+\frac{1}{4}},$ (15b)
holding for $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\phi)<\infty$, where $C^{\prime}_{q}>0$ is a constant. Estimate (15a) can be found in [8, (3.10)], while (15b) follows from (15a) and the Cauchy-Schwarz inequality, since we have $\nabla\phi\in L^{2}(\mathbb{R}^{4})$ and $\|\nabla\phi\|_{L^{2}(\mathbb{R}^{4})}\leq\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{1/4}$.
To see that any $v\in\mathcal{C}$ is bounded, we use the fact that $\nabla v\in L^{4}(\mathbb{R}^{4})$, which implies that $\mu(\{|\nabla v|>1\})<\infty$. Then, applying (15a) to the function $\psi=(|\nabla v|-1)^{+}$, we obtain that
$\gamma_{q}=\int_{|\nabla v|>1}|\nabla v|^{q}$ $\leq\int_{|\nabla v|>1}((|\nabla v|-1)^{+}+1)^{q}$
$\leq(\|\psi\|_{L^{q}(\mathbb{R}^{4})}+\mu(\{|\nabla v|>1\}))^{q}<\infty,$
for any $q\in[1,\infty)$. Consequently, it follows that the integrals
$\int_{B(x,1)}|\nabla v|^{q}$ $=\int_{\{|\nabla v|>1\}\cap B(x,1)}|\nabla v|^{q}+\int_{\{|\nabla v|\leq 1\}\cap B(x,1)}|\nabla v|^{q}$
$\leq\gamma_{q}+\mu(B(0,1)),$
are uniformly bounded on the unit balls of $\mathbb{R}^{4}$. A similar application of (15b) to $\phi=(|v|-1)^{+}$ shows that the integrals $\int_{B(x,1)}|v|^{q}$ are also uniformly bounded. Therefore the $W^{1,q}$ norm of $v$ is uniformly bounded on the unit balls, for any $q\in[1,\infty)$. This implies that $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, for any $\alpha\in(0,1)$.
A consequence of the uniform continuity of $v$ is that the set $\{|v|>\epsilon\}$ is bounded. Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint. Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction.
Now, let us assume the existence of a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ satisfying $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$. According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough). Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$. As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction. Therefore, we established that $\lim_{|x|\to\infty}v(x)=0$.
Finally, since the measures $\mu(\{|\nabla u_{j}|>1\})$ and $\mu(\{|u_{j}|>\varepsilon\})$ are uniformly bounded, the same arguments applied to the sequences $\psi_{j}=(|\nabla u_{j}|-1)^{+}$ and $\phi_{j}=(|u_{j}|-\epsilon)^{+}$ imply (14), and the fact that the $W^{1,q}$ norms of $\{u_{j}\}$ are uniformly bounded on the unit balls, for any $q\in[1,\infty)$. Thus, the sequence $\{u_{j}\}$ is equibounded and equicontinuous in $\mathbb{R}^{4}$. ∎
At this stage, we recall the following compactness result [13, Lemma 6].
@@
Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon\})\geq\delta>0$. Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/2\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/2\})\geq\beta$. Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/2\})\geq\beta$. On the other hand, the previous estimates imply that we can extract a subsequence still called $\{u_{j}\}$ such that
  1. 1.
    $u_{j}\to u$ uniformly on compact subsets of $\mathbb{R}^{4}$,
  2. 2.
@@
  4. 4.
    $\mu(B\cap\{|u|\geq\varepsilon/2\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
  5. 5.
    $\mu(\{|u|>\varepsilon\})\leq 1$,
  6. 6.
    $u\in\mathcal{C}$.
Properties 4. and 5. of the limit $u$ follow from the convergence in 1., which is established via the theorem of Ascoli and a diagonal argument. To check that $G(u)\in L^{1}$, we use assumption (5) and the inequality $G^{+}(v)\leq k|v|$, holding for $|v|\leq\sup_{j}\|u_{j}\|_{L^{\infty}(\mathbb{R}^{4})}$, and for a constant $k>0$. Then, (14) implies that the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$. Similarly, since we have $\int_{\mathbb{R}^{4}}G^{-}(u_{j})\leq\int_{\mathbb{R}^{4}}G^{+}(u_{j})$, the integrals $\int_{\mathbb{R}^{4}}G^{-}(u_{j})$ are uniformly bounded, and we obtain as previously that $\int_{\mathbb{R}^{4}}G^{-}(u)<\infty$. Therefore, $G(u)\in L^{1}$. To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$. Indeed, if $\eta\geq\epsilon$, this is clear in view of 5. above. Otherwise, if $0<\eta<\epsilon$, we have
$\mu(\{\eta<|u|\leq\epsilon\})\min\{|G(v)|:\eta\leq|v|\leq\epsilon\}\leq\int_{\eta\leq|u|\leq\epsilon}|G(u)|<\infty,$
thus $\mu(\{\eta<|u|\leq\epsilon\})<\infty\Rightarrow\mu(\{|u|>\eta\})<\infty$.
@@
$\mathcal{K}=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}|\ \phi\in C_{0}(\mathbb{R}^{4}),\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2}\right\},$ (16)
where $C_{0}$ is the space of continuous functions with compact support.
@@
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Then,
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,$ (17)
where $D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial^{2}\phi}{\partial x_{i}\partial x_{j}}$.

### v3

@@ 1 Introduction and main results
Due to the increasing interest for higher order models and their applications (e.g. in nonlinear elasticity [4] and mechanics [26]), it is also relevant to investigate ground state problems for the bilaplacian operator $\Delta^{2}$. This was done by Mederski and Siemienowski [23] for dimensions $d\geq 5$, and as an application they derived a biharmonic logarithmic Sobolev inequality (cf. [23, Theorem 1.4]).
The scope of the present paper is to study the limiting case of space dimension $4$. Our approach is based on the method introduced in [9], as well as on some tools borrowed from [23] (e.g. Pohozaev identity in [23, Theorem 1.2]). Following [23], we also extend the biharmonic logarithmic Sobolev inequality in [23, Theorem 1.4] to dimension $4$.
@@
where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
@@
$G(u_{0})>0\ \text{for some}\ u_{0},$ (6)
$\text{$\forall b>0,\ \forall\delta>0,\ \exists C_{\delta,b}>0$ such that }|g(u)|\leq C_{\delta,b}+\delta e^{b|u|^{4/3}},\quad\forall u\in\mathbb{R}^{4}.$ (7)
That is, the function $G$ has at $u=0$ a local (but not global) maximum equal to 0. We also point out that $G$ need not be differentiable at 0. For instance, we may have G(u)=|u|G(u)=-|u| near 0. On the other hand, it follows from (7) that
$\lim_{|u|\to\infty}\dfrac{|g(u)|}{e^{a|u|^{4/3}}}=0\ \text{and}\ \lim_{|u|\to\infty}\dfrac{|G(u)|}{e^{a|u|^{4/3}}}=0,\ \forall a>0.$ (8)
We shall see later that this growth condition is required to obtain compactness properties from the embedding of $W^{1,4}(\mathbb{R}^{4})$ in the Orlicz space $L_{A}(\mathbb{R}^{4})$, where the function $A$ is defined in (24).
The energy functional associated with system (2) is
@@
$\text{where }K(u)=\dfrac{1}{2}\int\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (10)
$\text{ and }V(u)=\int G(u(x))dx.$ (11)
@@
It is worth noting that $K(u)=\frac{1}{2}\int_{\mathbb{R}^{4}}|\Delta u|^{2}$, provided that $\nabla u\in D^{1,2}(\mathbb{R}^{4})$ (cf. Lemma 11). Thus, $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$.
@@
If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
@@
which clearly satisfies assumptions (3)-(7). Moreover, for this choice of $G$, one can see that $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$ (cf. Remark 6). Following [23], we derive in (14) below the biharmonic logarithmic Sobolev inequality in dimension $4$ (cf. [23, Theorem 1.4] for dimensions $d\geq 5$). It is a generalization for the bilaplacian operator of the classical logarithmic Sobolev inequality
@@
For any $u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m})$ such that $\int_{\mathbb{R}^{4}}|u|^{2}=1$, there holds
@@
with $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$, where $T>0$ is the minimum in (12). Moreover, the equality in (14) holds provided that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ and $\bar{u}$ is a ground state solution to (2). Conversely, if the equality in (14) holds for $u$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
@@
If $u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m})$, we have $(|u|^{2}\ln|u|)^{+}\in L^{1}(\mathbb{R}^{4})$ (cf. Remark 6), and hence the right hand side of (14) is well defined, although it could be -$\infty$ if $\int_{\mathbb{R}^{4}}(|u|^{2}\ln|u|)^{-}=+\infty$.
Remark 3.
Equality case in (14) is achieved for the normalized ground state solutions provided by Theorem 1.1. However, these solutions are not known explicitly, thus we neither know the exact value of the optimal constant in (14). By comparing it with the optimal constant (cf. [10]) of the classical logarithmic Sobolev inequality (13), we can at least obtain that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$.
Logarithmic Sobolev type inequalities appear in several equivalent formulations. The seminal inequality by Gross [18] was established in the Gaussian setting. Weissler’s approach [27] gives the Lebesgue analogue with sharp constant which under the unit normalization $\|u\|_{L^{2}(\mathbb{R}^{4})}=1$ leads to the form (13) (see also [20]).
In a series of papers [12, 13], Cotsiolis and Tavoularis proved sharp Sobolev type inequalities with explicit best constants for higher order fractional derivatives in the range $d>2s$. Most relevant here is their higher order fractional derivative logarithmic result [14], which for every $s>0$ and every $\alpha>0$ gives a sharp inequality with explicit optimal constants: for $u\in H^{s}(\mathbb{R}^{d})$,
$\int_{\mathbb{R}^{d}}|u|^{2}\ln\!\Big(\frac{|u|^{2}}{\|u\|_{2}^{2}}\Big)\,dx+\Big(d+\frac{d}{s}\ln\alpha+\ln\!\Big(s\,\frac{\Gamma(d/2)}{\Gamma(d/(2s))}\Big)\Big)\|u\|_{2}^{2}\\ \;\leq\;\frac{\alpha^{2}}{\pi^{s}}\|(-\Delta)^{s/2}u\|_{2}^{2}.$ (15)
In particular, for $s=1$ one recovers the Gross inequality (13) by setting $\|u\|_{2}=1$ and minimizing in $\alpha$.
Finally, Mederski and Siemianowski (for $d\geq 5$) derived a scalar biharmonic logarithmic Sobolev inequality tied to the bilaplacian, i.e. for $u\in D^{2,2}(\mathbb{R}^{d})$ with $\|u\|_{2}=1$,
$\int_{\mathbb{R}^{d}}|u|^{2}\ln|u|\,dx\;\leq\;\frac{d}{8}\ln\!\Big(C\int_{\mathbb{R}^{d}}|\Delta u|^{2}\,dx\Big),$ (16)
with some $0<C<\big(\frac{2}{\pi ed}\big)^{2}$. Moreover, they relate the equality to normalized ground state solutions of the associated biharmonic problem and note that the exact optimal constant is not explicitly computable. Note that, the case $s=2$ in (15) corresponds to the logarithmic inequality controlled by $\|\Delta u\|_{2}^{2}$, matching the inequality above.
The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in Section 4 we gather some technical lemmas.
@@
[…]
It is clear that the norms $\|D^{2}u_{j}\|_{L^{2}(\mathbb{R}^{4})}$ are uniformly bounded. Therefore, due to the Sobolev imbedding of $D^{1,2}(\mathbb{R}^{4})$ in $L^{4}(\mathbb{R}^{4})$ (cf. [28, Thm. 1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
Next, we show that $\{u_{j}\}\subset L^{q}_{loc}(\mathbb{R}^{4})$, $\forall q\in[1,+\infty)$, and estimate the $L^{q}$ norm of $u_{j}$.
@@
We have that
$\int_{|u_{j}|>\varepsilon/2}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (18)
@@
Our first claim is that given $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\,\phi)<\infty$, we have
$\|\phi\|_{L^{q}(\mathbb{R}^{4})}$ $\leq k_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\,\phi))^{\frac{1}{q}},\ \forall q\geq\frac{4}{3},$ (19)
and for a constant $k_{q}>0$. To establish (19), we follow the steps of the proof of [9, inequality (3.10)]. Indeed, it holds (cf. [25, Theorem p. 125]):
$\|\phi\|_{L^{\frac{4}{3}}(\mathbb{R}^{4})}$ $\leq k\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})},\ \forall\phi\in C^{1}_{0}(\mathbb{R}^{4}),$ (20)
for a constant $k>0$. Next, by applying (20) to the functions $\phi^{n}$, $n=2,3,\ldots$, we get by induction
$\|\phi\|_{L^{\frac{4n}{3}}(\mathbb{R}^{4})}$ $\leq k_{n}\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})}^{\frac{1}{n}}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}^{\frac{n-1}{n}},$ (21)
for a constant $k_{n}>0$, and by interpolation
$\|\phi\|_{L^{q}(\mathbb{R}^{4})}$ $\leq k_{q}\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})}^{\frac{4}{3q}}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}^{1-\frac{4}{3q}},\ \forall q\geq\frac{4}{3},$ (22)
for a constant $k_{q}>0$. Moreover, smoothing by convolution, we notice that (22) still holds for functions $\phi\in L^{\infty}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $\phi$ has compact support. On the other hand, a standard cut-off and truncation argument shows the validity of (22) for functions $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $\mu(supp\,\phi)<\infty$. Finally, (19) follows from (22) by Hölder’s inequality.
To conclude the proof of the lemma, we apply (19) to $\phi=(|u_{j}|-\varepsilon/2)^{+}$, and obtain
$\|(|u_{j}|-\varepsilon/2)^{+}\|_{L^{q}(\mathbb{R}^{4})}\leq k_{q}\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})},$
which implies (18). ∎
At this stage, we recall the following compactness result [19, Lemma 6].
@@
Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon/2\})\geq\delta>0$. Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/4\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/4\})\geq\beta$. Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/4\})\geq\beta$. On the other hand, Lemma 1 implies that given a bounded set $Q\subset\mathbb{R}^{4}$, the norms $\|u_{j}\|_{L^{q}(Q)}$ are uniformly bounded. Consequently, we can extract a subsequence still called $\{u_{j}\}$ such that
  1. (i)
    $u_{j}\to u$ in $L^{q}(Q)$, $\forall q\in[1,+\infty)$, $\forall Q\subset\mathbb{R}^{4}$ bounded,
  2. (ii)
    $u_{j}\to u$ a.e. in $\mathbb{R}^{4}$,
  3. (iii)
@@
  5. (v)
    $\mu(B\cap\{|u|\geq\varepsilon/4\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
  6. (vi)
    $\mu(\{|u|>\varepsilon/2\})\leq 1$,
  7. (vii)
    $u\in\mathcal{C}$,
Properties (v) and (vi) of the limit $u$ follow from the a.e. convergence in (ii). To check that $G(u)\in L^{1}$, we use assumption (8) and the embedding
$W^{1,4}(\mathbb{R}^{4})\hookrightarrow L_{A}(\mathbb{R}^{4}),$ (23)
where $L_{A}$ is the Orlicz space corresponding to the function
$A(t)=\exp(t^{4/3})-1-t^{4/3}-\dfrac{1}{2}t^{8/3}$ (24)
(cf. Trudinger’s theorem in [1, §8.29]). Since the sequence $\{(|u_{j}|-\epsilon/2)^{+}\}$ is uniformly bounded in $W^{1,4}(\mathbb{R}^{4})$, it follows that there exists $k_{1}>0$, such that the integrals $\int_{\mathbb{R}^{4}}A(k_{1}(|u_{j}|-\epsilon/2)^{+})$ are uniformly bounded. Moreover, it holds $G^{+}(u_{j})\leq k_{2}A(k_{1}(|u_{j}|-\epsilon/2)^{+})$, for a constant $k_{2}>0$ (cf. (5) and (8)). Consequently, the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$. Similarly, since we have $\int_{\mathbb{R}^{4}}G^{-}(u_{j})\leq\int_{\mathbb{R}^{4}}G^{+}(u_{j})$, the integrals $\int_{\mathbb{R}^{4}}G^{-}(u_{j})$ are uniformly bounded, and we obtain as previously that $\int_{\mathbb{R}^{4}}G^{-}(u)<\infty$. Therefore, $G(u)\in L^{1}$. To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\alpha\})<\infty$, $\forall\alpha>0$. Indeed, if $\alpha\geq\epsilon/2$, this is clear in view of (vi) above. Otherwise, if $0<\alpha<\epsilon/2$, we have in view of hypothesis (5):
$\mu(\{\alpha<|u|\leq\epsilon/2\})\min\{|G(v)|:\alpha\leq|v|\leq\epsilon/2\}\leq\int_{\alpha\leq|u|\leq\epsilon/2}|G(u)|<\infty,$
thus $\mu(\{\alpha<|u|\leq\epsilon/2\})<\infty\Rightarrow\mu(\{|u|>\alpha\})<\infty$. This completes the proof of property (vii) above.
@@
$\mathcal{K}:=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}\ |\ \phi\in L^{1}_{loc},\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2},\ \mu(supp\,\phi)<\infty\right\}.$ (25)
Remark 4.
Due to (19) we have that $\mathcal{K}\subset L^{q}(\mathbb{R}^{4})$, $\forall q\in[1,\infty)$.
@@
Let $\phi\in\mathcal{K}$ be such that
$\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0.$ (26)
Then,
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,\text{ where }D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\partial^{2}_{ij}u\cdot\partial^{2}_{ij}\phi.$ (27)
Note.
We shall see in the proof of Lemma 3 that $G(u+\phi)\in L^{1}(\mathbb{R}^{4})$, $\forall\phi\in\mathcal{K}$.